A bookmark has an area of 18 square centimeters and a perimeter of 22 centimeters. What are the dimensions of the bookmark?
Answer by Alan3354(69443)

A bookmark has an area of 18 square centimeters and a perimeter of 22 centimeters. What are the dimensions of the bookmark?
---------
2L + 2W = 22
L + W = 11
L*W = 18
-------
Find a pair of factors of 18 what add to 11.